what is the solution of 2/x-3/y=2;3/x+4/y=25?
Answer by kollanag(3)

2/x-3/y = 2
2y-3x = 2xy
3/x+ 4/y = 25
3y + 4x = 25xy
y = xy + 3x/2
y = x(y+3/2)
y = x(25y/3-4/3)
y + 3/2 = 25y/3 - 4/3
2y + 3 = 50y/3 - 8/3
6y + 9 = 5oy - 8
44y = 17
y = 17/44
and
2/x - 3* 44/17 = 2
2/x = 2 + 132/17
2/x = 166/17
1/x = 83/17
so x = 17/83 and y = 17/44